40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 272     Entry time: Thu Aug 12 00:10:44 2010
Author: Frank 
Type: Summary 
Category: RefCav 
Subject: both cavities locked ~24h 

this afternoon both cavities were locked since 24h. The initial problem was to find the right temperature for one of the cavities if we keep the other one fixed at 35 degrees. So now the reference cavity is held at 35 degrees, (because the time constant it needs to settle is much higher), while the analyzer cavity is at 35.3 degrees. With those setting the VCO input signal is almost at zero, having about +/-10MHz of tuning range.

ELOG V3.1.3-